OpenStudy (anonymous):

Use substitution. 2x + 3y = 5 4x - y = 17 Use linear combinations: 2x + 3y = 5 4x - y = 17

OpenStudy (anonymous):

1) 2x + 3y = 5 4x - y = 17 First, isolate one of the variables. 4x - y =17 4x = 17 + y y = 4x - 17 Then, put what you got for y into the other equation. 2x + 3(4x-17)=5 2x + 12x - 51 = 5 14x - 51 = 5 14x = 56 x = 56/14 x = 4 THEN, plug your answer for x back into one of the equations, and solve for y! 2(4) - 3y = 5